OpenStudy (anonymous):

stretch the asymptotes and graph the function y=5/x-2+3

OpenStudy (mathmale):

Please use parentheses to indicate which math operations are to be performed first. Did you mean 5 / (x-2) or did you mean 5 / x ? Also, what do you mean by "stretch the asymptotes?"

OpenStudy (anonymous):

5/(x-2)

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

OK. We have y = 5 / (x+2) +3. First: How does one locate asymptotes?

OpenStudy (anonymous):

I'm really lost with this. Can you break it down for me?

OpenStudy (mathmale):

Let's build upon what you already know. What is an "asymptote?" Where have you seen asymptotes before?

OpenStudy (anonymous):

i just saw them once in a lesson

OpenStudy (anonymous):

they werent really explained

OpenStudy (mathmale):

OK. We have y = 5 / (x+2) +3 Next time, please ask questions about whatever you found to be poorly explained. That way someone could help you more efficiently. In the above equation, what would happen if x = - 2?

OpenStudy (anonymous):

2?

OpenStudy (mathmale):

Another way in which to ask this same question: Can we let x=-2 in the given equation? Try it and share your results.

OpenStudy (anonymous):

y = 5 / (-2+2) +3

OpenStudy (mathmale):

That's right. Can you take the math any further than that?

OpenStudy (anonymous):

y = 5 / (-2+2) +3 y = 5 /+3

OpenStudy (mathmale):

Actually, it'd be\[y=\frac{ 5 }{ 0 }+3\] Make a judgment. Is this OK or not OK? Why?

OpenStudy (anonymous):

idk what would make it not ok or ok.

OpenStudy (anonymous):

y = 5 / (-2+2) +3 y = 5/0 +3

OpenStudy (mathmale):

Have you heard the phrase, "division by zero is .... "?

OpenStudy (anonymous):

I havent

OpenStudy (mathmale):

Division by zero is undefined. We can't divide by zero, period. This limitation means that you cannot let x=2. Why? Because then the denominator of 5 / (x-2) would be zero, which is not permitted. We'll need to come back to this later. But for now, having identified the x value for which the function is undefined, we can write, "the vertical asymptote of the graph of this function is the vertical line x=2. You could draw a set of coordinate

OpenStudy (mathmale):

axes and then draw the vertical line x=2 as well. Your graph can NOT cross this vertical axis.

OpenStudy (mathmale):



OpenStudy (mathmale):

Once again, your graph can NOT cross the vertical line x=2, which is your "vertical asymptote."

OpenStudy (anonymous):

so that is when lines to not cross?

OpenStudy (anonymous):

its kind of coming back to me now i thought it was when lines dont cross the axis

OpenStudy (mathmale):

I'm sorry, Dani, but I need to get off the Internet now. But I'd be glad to continue this discussion. Tag me later on today, or repost this question. To answer you question: It's the graph that can't cross the verticql asymptote, x=2. I will go into that further when we reconnect. For now, I challenge you to look at the equation given and try to figure out what happens to y as x becomes larger and larger. This is how you determine your "horiz. asymptote."